OpenStudy (anonymous):

solve the formula r=√2a/θ for θ

OpenStudy (saifoo.khan):

\[\Huge r = \sqrt{\frac{2a}{\theta}}\]

OpenStudy (anonymous):

θ=√2a/r

OpenStudy (anonymous):

saifoo , what you changed then?

OpenStudy (saifoo.khan):

i didn't changed anything i was asking the right format of the question!!

OpenStudy (anonymous):

yes thats the format

OpenStudy (anonymous):

oh, your question was not clear, it got no ?

OpenStudy (anonymous):

i just needed help on solving it for theta

OpenStudy (saifoo.khan):

mine?